Spring Break is coming up and you are headed to Vegas! You decide that you are going to play roulette and want to know the probability of landing on certain numbers. The standard American roulette wheel has a total of 38 pockets. There are two green pockets, 18 red pockets, and 18 black pockets. With each spin of the wheel, the ball has an equal probability of landing in any pocket.

1a. With any given spin of the wheel, what is the probability that you land in a green pocket? Give your answer as a proportion and round to 4 decimal places. Show your work and use correct notation.

1b. Convert the probability from question 1a to a percentage. Show your work.

2a. With any given spin of the wheel, what is the probability that you land in a black OR red pocket? Give your answer as a proportion and round to 4 decimal places. Show your work and use correct notation.

2b. Convert the probability from question 2a to a percentage. Show your work.
Answer #1

Answer

total number of pockets =38

number of green pockets = 2

number of black pockets = 18

number of red pockets = 18

(1a) We know that probability =(favorable outcome)/(total outcome)

here total outcome = 38 and favorable outcome = 2 for green pockets

So, we get

Probability = (2/38)= 0.0526

(1b) Converting probability to percentage, we need to multiply the probability by 100.

So, we get

Percentage = probability*100 = 0.0526*100 = 5.26%

(2a) We know that probability =(favorable outcome)/(total outcome)

here total outcome = 38 and favorable outcome = 18 for black pockets and 18 for red pockets

So, we get

Probability = (18/38)+ (18+38)= 0.4737 + 0.4737 = 0.9474

(2b) Converting probability to percentage, we need to multiply the probability by 100.

So, we get

Percentage = probability*100 = 0.9474*100 = 94.74%